# ProxNF: Neural Field Proximal Training for High-Resolution 4D Dynamic Image Reconstruction

### v1

@@ II-B Neural fields
A neural field (NF) is a neural network that provides a representation of continuous objects, such as images, videos, and audio signals [20]. An NF can be described as a neural network $\Phi_{\boldsymbol{\xi}}:\Omega_{T}\rightarrow\mathbb{R}$ that acts as a map between spatiotemporal coordinates to the values of an image at those coordinates. NF architectures represent a finite-dimensional, nonlinear topological subspace of $L^{2}(\Omega_{T})$ parametrized by a set of network parameters $\boldsymbol{\xi}\in\mathbb{R}^{N_{p}}$. For many classes of images, in particular biomedical images[23, 29], NF can provide accurate representations using a fixed network architecture and requiring a relatively small number of parameters $N_{p}$ compared to the number of voxels needed to represent the image with a comparable resolution [17]. NFs are then able to significantly reduce the memory requires for storing high-resolution, high-dimensional images.
@@ II-B1 Spatiotemporal image reconstruction using neural fields
Advantages of using an NF for spatiotemporal image reconstruction in Eq. (12) include reduced memory requirements for optimization. However, this does come with a few disadvantages. First, optimizing over the space of network parameters to train the NF is a nonconvex optimization problem, even in the case when the original problem in Eq. (4) is convex. The nonconvex nature of this problem means there are no guarantees on finding a global minimizer; additionally non-convex problems are in general more difficult to solve than convex problems. Second, while computationally efficient implementations of several canonical discrete-to-discrete imaging operators are readily available, the evaluation of a continuous-to-discrete imaging operator acting on the NF representation, $\mathcal{H}(\Phi_{\boldsymbol{\xi}})$ is computationally expensive and further slows down training.
@@ II-B2 Partition of unity networks
One particular NF network architecture is the partition of unity network (POUnet) [30], which has previously been utilized for spatiotemporal image reconstruction in two spatial dimensions [13]. The POUnet consists of three components: a trainable partition of unity $\boldsymbol{\Psi}_{\boldsymbol{\eta}}:\Omega_{T}\rightarrow[0,1]^{P}$ parameterized by a set of weights $\boldsymbol{\eta}\in\mathbb{R}^{W}$, a trainable coefficient matrix $\boldsymbol{C}\in\mathbb{R}^{L\times P}$, and a fixed set of spatiotemporal basis functions $\boldsymbol{B}:\Omega_{T}\rightarrow\mathbb{R}^{L}$. Specifically, the POUnet NF representation reads
$\Phi_{(\boldsymbol{\eta},\boldsymbol{C})}(\mathbf{x})=\boldsymbol{\Psi}_{\boldsymbol{\eta}}^{T}(\mathbf{x})\boldsymbol{C}^{T}\boldsymbol{B}(\mathbf{x}),$
where the partition of unity network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}(\mathbf{x})=(\psi_{\boldsymbol{\eta}}^{1}(\mathbf{x}),\ldots,\psi_{\boldsymbol{\eta}}^{P}(\mathbf{x}))$ is such that
$\sum_{p=1}^{P}\psi^{p}_{\boldsymbol{\eta}}(\mathbf{x})=1,\text{ and }\psi^{p}_{\boldsymbol{\eta}}(\mathbf{x})\geq 0.$
@@
The POUnet architecture is advantageous because of this fast optimization strategy, particularly in the case of convex cost functionals $\mathcal{J}$ in which updates of the coefficients $\boldsymbol{C}$ are also convex. Furthermore, the POUnet architecture can provide theoretical accuracy guarantees based on a priori error estimates for piecewise polynomial interpolation [30].
@@ III-A ProxNF: a proximal training method for neural field dynamic image reconstruction
Previously, it has been shown that NFs can be utilized for small-scale spatiotemporal image reconstruction problems and trained with gradient based methods[13, 14]. However, these gradient-based methods become very slow when scaling to high-resolution, high-dimensional problems. This lack of scalability is due to the difficulties outlined in Section II-B1 of nonconvexity and expensive imaging operator evaluations.
Notably, these disadvantages are greatly reduced in the supervised embedding problem in Eq. (11) for a few reasons. First although this problem is nonconvex, embedding is very well studied and great amounts of effort have been made to design gradient-based training methods that can approximately find a “good” local minimum. Here a “good” local minimum refers to a local minimum with a sufficiently low suboptimality gap. Second, evaluating the embedding training objective and its gradient require no calls to the imaging operator and is computationally inexpensive. This means that the supervised problem can be solved relatively quickly and easily when compared to Eq. (12).
To leverage the existing capabilities for solving NF embedding problems, this work proposes a novel formulation of the spatiotemporal image reconstruction problem as constrained minimization problem, in which the sought-after object estimate $f^{*}$ is constrained to the manifold $\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}\coloneqq\{f\in L^{2}(\Omega_{T})|f={\Phi}_{\xi},\ \xi\in\mathbb{R}^{W}\}$ of all functions that can be represented exactly by the chosen NF architecture. That is, the object estimate $f^{*}$ is defined as
$f^{*}\coloneqq\operatorname*{argmin}_{f\in\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}}\frac{1}{2\sigma^{2}}\left\|\mathcal{H}(f)-\boldsymbol{d}\right\|^{2}+\mathcal{R}(f).$ (14)
In this work, a stochastic generalized plug-and-play method, referred to as ProxNF, is developed to approximately solve Eqn. (14), as outlined in Algorithm 1. Plug-and-play methods [31, 32, 33, 34] are a generalization of the proximal gradient algorithm [35, 36] in which the proximal operator is replaced by some denoising method. To this aim, a generalized proximal operator (denoiser) that computes a regularized projection of a spacetime object estimate $f\in L^{2}(\Omega_{T})$ onto the manifold $\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}$ is introduced. Its evaluation requires the solution of a regularized embedding problem defined as
$\operatorname{Prox}_{\alpha,\mathcal{R}}^{\textnormal{NF}}(f):=\operatorname*{argmin}_{\tilde{f}\in\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}}\frac{1}{2\alpha}\|\tilde{f}-f\|^{2}+\mathcal{R}(\tilde{f}).$ (15)
@@
$g^{n}=\nabla_{f}\mathcal{L}(f^{n}),\ f^{n+1}=\operatorname{Prox}_{\alpha^{n},\mathcal{R}}^{\textnormal{NF}}(f^{n}-\alpha^{n}g^{n}),$ (16)
@@
This method then reframes the original self-supervised problem as a series of, easier to solve, supervised subproblems. In this work the update direction $g$ is chosen to be the gradient of the data-fidelity term, but could be generalized to include other choice of update direction (potentially conjugate gradient or (Quasi-)Newton directions).
Approximating the continuous-to-discrete the imaging operator $\mathcal{H}(f)$ with $\boldsymbol{H}(\Pi_{M,K}f)$, the update direction $g$ in Eq. (16) is numerically computed as
$g^{n}=\Pi_{M,K}^{*}\nabla_{\mathbf{F}}\mathbf{L}(\Pi_{M,K}f^{n}),$
@@
where $\Pi_{M,K}^{*}:\mathbb{R}^{M\times K}\rightarrow\mathbb{L}^{2}(\Omega_{T})$ is the adjoint of $\Pi_{M,K}$ defined in Eqn (7). However, explicitly forming this update direction $g^{n}$ requires to store an expansion coefficient matrix $\mathbf{G}^{n}\in\mathbb{R}^{M\times K}$ and thus can compromise the memory efficiency provided by the NF representation of the object. Memory efficiency can be preserved by computing an update direction $\tilde{g}$ using stochastically subsampled frames. For a stochastically selected collection of frames $\{k_{1},\ldots,k_{J}\}$ (with $J\ll K$), the update direction is then computed as
@@
$\tilde{g}^{n}=\Pi_{M,K}^{*}\sum_{j=1}^{J}\nabla_{\mathbf{f}_{k_{j}}}\mathbf{L}_{k_{j}}(\Pi_{M,K}^{k_{j}}f^{n})\otimes\mathbf{e}_{k_{j}},$
@@
which only requires storing an expansion coefficient matrix $\tilde{\mathbf{G}}^{n}\in\mathbb{R}^{M\times J}$. This proposed method is outlined in Algorithm 1.
It is worth noting that Nesterov acceleration is not used Algorithm 1. This is because proximal splitting methods combining randomized evaluation of the data-fidelity gradient with Nesterov acceleration lack theoretical convergence guarantees and may be susceptible to instabilities due to accumulation of errors in the momentum term [37, 38, 39, 40].
@@
             Compute frame-wise coefficients
@@
Compute frame-wise directions
@@
      Form the update direction
$\tilde{g}^{n}=\Pi_{M,K}^{*}\sum_{j=1}^{J}\mathbf{g}_{k_{j}}^{n}\otimes\mathbf{e}_{k_{j}}.$
For a step size $\alpha$, apply proximal update
$f^{n+1}\leftarrow\operatorname{Prox}^{\textnormal{NF}}_{\alpha,\mathcal{R}}(f^{n}-\alpha\tilde{g}^{n})$
@@
Algorithm 1 ProxNF: Dynamic image reconstruction with neural fields via proximal splitting
III-B Adapted partition of unity neural fields
This work proposes utilizing modified partition of unity network of the form
$\Phi_{(\boldsymbol{\eta},\boldsymbol{C})})(\boldsymbol{r},t)=\boldsymbol{\Psi}^{T}_{\boldsymbol{\eta}}(t)^{T}\boldsymbol{C}^{T}\boldsymbol{B}(\boldsymbol{r}),$
where the partition of unity network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ only depends on the time coordinate $t$ and the basis functions $\boldsymbol{B}$ only depends on the spatial coordinate $\boldsymbol{r}$. In particular, $\boldsymbol{B}(\boldsymbol{r})=\begin{pmatrix}\beta_{1}(\boldsymbol{r})&\ldots&\beta_{M}(\boldsymbol{r})\end{pmatrix}$ shares the same basis functions $\{\beta_{m}\}_{m=1}^{M}$ in Eqn. (5). This representation allows for the high spatial resolution provided by traditional discretization methods and smooth temporal transitions while still exploiting spatiotemporal redundancies. Moreover, this architecture greatly simplifies discrete projection operations. That is for a selected frame $k$,

### v3

@@ II-B Neural fields
A neural field (NF) is a neural network that provides a representation of continuous objects, such as images, videos, and audio signals [20]. An NF can be described as a neural network $\Phi_{\boldsymbol{\xi}}(\boldsymbol{x}):\Omega_{T}\rightarrow\mathbb{R}$ that maps every spatiotemporal point $\boldsymbol{x}=(\boldsymbol{r},t)\in\Omega_{T}$ to the values of an image at that point. NF architectures represent a finite-dimensional, nonlinear topological subspace of $L^{2}(\Omega_{T})$ parametrized by a set of network parameters $\boldsymbol{\xi}\in\mathbb{R}^{N_{p}}$. For many classes of images, in particular biomedical images[23, 36, 37], NFs can provide accurate representations using a fixed network architecture and requiring a relatively small number of parameters $N_{p}$ compared to the number of voxels needed to represent the image with a comparable resolution [17]. NFs are then able to significantly reduce the memory required for storing high-resolution, high-dimensional images.
@@ II-B1 Spatiotemporal image reconstruction using neural fields
Advantages of using an NF for spatiotemporal image reconstruction in Eq. (14) include reduced memory requirements for optimization. However, this does come with a few disadvantages. First, optimizing over the space of network parameters to train the NF is a nonconvex optimization problem, even in the case when the original problem in Eq. (4) is convex. The nonconvex nature of this problem means there are no guarantees on finding a global minimizer; additionally, non-convex problems are in general more difficult to solve than convex problems. Second, while computationally efficient implementations of several canonical discrete-to-discrete imaging operators are readily available, the evaluation of a continuous-to-discrete imaging operator acting on the NF representation, $\mathcal{H}(\Phi_{\boldsymbol{\xi}})$ is computationally expensive and further slows down training.
@@ II-B2 Partition of unity networks
One particular NF network architecture is the partition of unity network (POUnet) [38], which has previously been utilized for spatiotemporal image reconstruction in two spatial dimensions [13]. The POUnet consists of three components: a trainable partition of unity $\boldsymbol{\Psi}_{\boldsymbol{\eta}}:\Omega_{T}\rightarrow[0,1]^{P}$ parameterized by a set of weights $\boldsymbol{\eta}\in\mathbb{R}^{W}$, a trainable coefficient matrix $\boldsymbol{C}\in\mathbb{R}^{L\times P}$, and a fixed set of spatiotemporal basis functions $\boldsymbol{B}:\Omega_{T}\rightarrow\mathbb{R}^{L}$. Specifically, the POUnet NF representation is given by
$\Phi_{(\boldsymbol{\eta},\boldsymbol{C})}(\boldsymbol{x})=\boldsymbol{\Psi}_{\boldsymbol{\eta}}^{T}(\boldsymbol{x})\boldsymbol{C}^{T}\boldsymbol{B}(\boldsymbol{x}),\quad\forall\boldsymbol{x}=(\boldsymbol{r},t)\in\Omega_{T},$ (15)
where the partition of unity network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}(\boldsymbol{x})=(\psi_{\boldsymbol{\eta}}^{1}(\boldsymbol{x}),\ldots,\psi_{\boldsymbol{\eta}}^{P}(\boldsymbol{x}))$ is such that
$\sum_{p=1}^{P}\psi^{p}_{\boldsymbol{\eta}}(\boldsymbol{x})=1,\text{ and }\psi^{p}_{\boldsymbol{\eta}}(\boldsymbol{x})\geq 0.$
@@
The advantage of using the above alternating optimization strategy is that, due to the linearity of $\Phi_{(\boldsymbol{\eta},\boldsymbol{C})}$ with respect to $\boldsymbol{C}$, the first minimization problem admits a unique global minimum when $\mathcal{J}$ is strongly convex. Thanks to this property, the POUnet representation possesses theoretically guaranteed approximation properties based on a priori error estimates for piecewise polynomial interpolation [38].
@@ III-A ProxNF: a proximal training method for neural field dynamic image reconstruction
Previously, it has been shown that NFs can be utilized for small-scale spatiotemporal image reconstruction problems and trained with gradient-based methods[13, 14]. However, these gradient-based methods become very slow when scaling to high-resolution, high-dimensional problems. This lack of scalability is due to the difficulties outlined in Section II-B1 of nonconvexity and expensive imaging operator evaluations.
Notably, these disadvantages are greatly reduced in the supervised embedding problem in Eq. (13) for a few reasons. First, although this problem is nonconvex, embedding is very well studied and great amounts of effort have been made to design gradient-based training methods that can approximately find a “good” local minimum. Here a “good” local minimum refers to a local minimum with a sufficiently low suboptimality gap. Second, evaluating the embedding training objective and its gradient requires no calls to the imaging operator and is computationally inexpensive. This means that the supervised problem can be solved relatively quickly and easily when compared to Eq. (14).
III-A1 Continuous formulation
To leverage the existing capabilities for solving NF embedding problems, this work proposes a novel formulation of the spatiotemporal image reconstruction problem as a constrained minimization problem, in which the sought-after object estimate $f^{*}$ is constrained to the manifold $\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}\coloneqq\{f\in\mathbb{U}|f={\Phi}_{\xi},\ \xi\in\mathbb{R}^{W}\}$ of all square integrable spatiotemporal functions that can be represented exactly by the chosen NF architecture. That is, the object estimate $f^{*}$ is defined as
$f^{*}\coloneqq\operatorname*{argmin}_{f\in\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}}\sum_{k=1}^{K}\mathcal{L}_{k}(f(\cdot,t_{k}))+\mathcal{R}(f),$ (17)
where, recalling Eq. (5), the data fidelity term is given by
$\begin{split}\sum_{k=1}^{K}\mathcal{L}_{k}(f(\cdot,t_{k}))&=\sum_{k=1}^{K}\frac{1}{2\sigma^{2}}\left\|\mathcal{H}_{k}(f(\cdot,t_{k})-\boldsymbol{d}_{k}\right\|^{2}\\ {}&=\frac{1}{2\sigma^{2}}\left\|\mathcal{H}(f)-\boldsymbol{d}\right\|^{2}.\end{split}$
In this work, a stochastic generalized plug-and-play method, referred to as ProxNF, is developed to approximately solve Eqn. (17), as summarized in Algorithm 1. Plug-and-play (PnP) methods [39, 40, 41, 42] are a generalization of the proximal gradient algorithm [43, 44] in which the proximal operator is replaced by some denoising method. To this aim, a generalized proximal operator (denoiser) that computes a regularized projection of a spacetime object estimate $f\in\mathbb{U}$ onto the manifold $\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}$ is introduced. Its evaluation requires the solution of a regularized embedding problem defined as
$\operatorname{Prox}_{\alpha\mathcal{R}}^{\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}}(f):=\operatorname*{argmin}_{\tilde{f}\in\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}}\frac{1}{2\alpha}\|\tilde{f}-f\|_{2}^{2}+\mathcal{R}(\tilde{f}).$ (18)
@@ III-A1 Continuous formulation
$g^{n}=\nabla_{f}\mathcal{L}(f^{n}),\ f^{n+1}=\operatorname{Prox}_{\alpha^{n}\mathcal{R}}^{\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}}(f^{n}-\alpha^{n}g^{n}),$ (19)
@@
This method then reframes the original self-supervised problem as a series of, easier to solve, supervised subproblems. In this work, the update direction $g$ is chosen to be the gradient of the data-fidelity term but could be generalized to include other choice of update direction (potentially conjugate gradient or (Quasi-)Newton directions).
III-A2 Discretized formulation
Approximating the continuous-to-discrete imaging operator $\mathcal{H}(\cdot)$ with $\mathcal{H}_{\Pi}(\cdot):=\mathbf{H}(\Pi_{M,K}\cdot)$, the discretized counterpart of the minimization problem in Eqn. (17) reads
$f_{\Pi}^{*}\coloneqq\operatorname*{argmin}_{f\in\mathcal{M}_{{\Phi}_{\boldsymbol{\xi}}}}\sum_{k=1}^{K}\boldsymbol{L}_{k}(\Pi^{k}_{M,K}f)+\mathcal{R}(f),$ (20)
where the data fidelity term is given by
$\begin{split}\sum_{k=1}^{K}\boldsymbol{L}_{k}(\Pi^{k}_{M,K}f)&=\sum_{k=1}^{K}\frac{1}{2\sigma^{2}}\left\|\mathbf{H}_{k}(\Pi^{k}_{M,K}f)-\boldsymbol{d}_{k}\right\|^{2}\\ {}&=\frac{1}{2\sigma^{2}}\left\|\mathbf{H}(\Pi_{M,K}f)-\boldsymbol{d}\right\|^{2}.\end{split}$
Above and in what follows, the subscript $\Pi$ in $f_{\Pi}^{*}$ is used to highlight the approximation introduced by replacing the continuous-to-discrete imaging operator $\mathcal{H}(\cdot)$ with the discretized one $\mathcal{H}_{\Pi}(\cdot)$.
The update direction $g_{\Pi}\in\mathbb{U}_{M,K}$, i.e. the counterpart of $g$ in Eq. (19) for the discretized problem in Eq. (20), is numerically computed as
$\begin{split}g_{\Pi}^{n}&=\Pi_{M,K}^{*}\nabla_{\mathbf{F}}\mathbf{L}(\Pi_{M,K}f^{n})\\ {}&=\frac{1}{\sigma^{2}}\Pi_{M,K}^{*}\mathbf{H}^{t}\left(\mathbf{H}(\Pi_{M,K}f)-\boldsymbol{d}\right),\end{split}$ (21)
@@ III-A2 Discretized formulation
where $\Pi_{M,K}^{*}:\mathbb{R}^{M\times K}\rightarrow\mathbb{L}^{2}(\Omega_{T})$ is the adjoint of $\Pi_{M,K}$ defined in Eqn (8). However, explicitly forming the update direction $g_{\Pi}^{n}\in\mathbb{U}_{M,K}$ in Eqn. (21) requires to store an expansion coefficient matrix $\mathbf{G}^{n}:=\nabla_{\mathbf{F}}\mathbf{L}(\Pi_{M,K}f^{n})\in\mathbb{R}^{M\times K}$ and thus can compromise the memory efficiency provided by the NF representation of the object. Memory efficiency can be preserved by computing an update direction $\tilde{g}_{\Pi}$ using stochastically subsampled frames. For a stochastically selected collection of frames $\{k_{1},\ldots,k_{J}\}$ (with $J\ll K$), the update direction is then computed as
@@
$\begin{split}\tilde{g}_{\Pi}^{n}&=\Pi_{M,K}^{*}\frac{K}{J}\sum_{j=1}^{J}\nabla_{\mathbf{f}_{k_{j}}}\mathbf{L}_{k_{j}}(\Pi_{M,K}^{k_{j}}f^{n})\otimes\mathbf{e}_{k_{j}}\\ &=\frac{K}{J}\sum_{j=1}^{J}\left(\Pi_{M,K}^{k_{j}}\right)^{*}\nabla_{\mathbf{f}_{k_{j}}}\mathbf{L}_{k_{j}}(\Pi_{M,K}^{k_{j}}f^{n})\\ {}&=\frac{K}{J}\sum_{j=1}^{J}\left(\Pi_{M,K}^{k_{j}}\right)^{*}\boldsymbol{g}_{k_{j}}^{n},\end{split}$ (22)
with
$\boldsymbol{g}_{k_{j}}^{n}=\nabla_{\mathbf{f}_{k_{j}}}\mathbf{L}_{k_{j}}(\Pi_{M,K}^{k_{j}}f^{n})=\mathbf{H}_{k_{j}}^{t}\left(\mathbf{H}_{k_{j}}\Pi_{M,K}^{k_{j}}f^{n}-\boldsymbol{d}_{k}\right).$
@@
Representing $\tilde{g}_{\Pi}^{n}$ only requires storing an expansion coefficient matrix $\tilde{\mathbf{G}}^{n}\in\mathbb{R}^{M\times J}$, whose $j$-th column is $\boldsymbol{g}_{k_{j}}^{n}$.
This proposed method is summarized in Algorithm 1, where the subscript $\Pi$ is omitted for ease of notation. Note that in the algorithm the spatiotemporal function $f^{n}$ ($n\geq 0$) is represented using an NF and only the network parameters $\boldsymbol{\xi}^{n}\in\mathbb{R}^{N_{p}}$ need to be stored. Similarly, the spatiotemporal function $\tilde{g}^{n}$ ($n\geq 0$) stemming from the randomly subsampled gradient of the data fidelity term is represented by the coefficients of the dense matrix $\tilde{\mathbf{G}}^{n}\in\mathbb{R}^{M\times J}$ with $J\ll K$. Thus, Algorithm 1 can be implemented in a memory-efficient manner, without the need to store an explicit representation of the spatiotemporal object estimate as an $\mathbb{R}^{M\times K}$ coefficient matrix.
Remark 1.
Algorithm 1 finds an approximate solution to the minimization problem in Eqn. (20) by use of proximal gradient descent (i.e., the so-called forward-backward splitting). In principle, also the alternating direction method of multipliers (ADMM, [32]) or backward-backward splitting (e.g. Douglas-Rachford algorithm, [45, 46]) could be employed to find an approximate solution of Eqn. (20). However, it is worth to note that while Algorithm 1 allows for a memory-efficient implementation that only requires storing the NF parameters $\boldsymbol{\xi}$ and the vectors $\boldsymbol{g}^{n}_{k_{j}}$, this is not possible using ADMM. In fact, updating the Lagrangian multiplier in the ADMM method would require the solution of an extra embedding step.
@@
             Render the NF at the $k_{j}$-th frame
@@
            Compute the gradient of the discretized data fidelity term w.r.t. the $k_{j}$-th frame coefficients
@@
      Form the update direction according to Eqn. (22)
$\tilde{g}^{n}=\frac{K}{J}\sum_{j=1}^{J}\left(\Pi_{M,K}^{k_{j}}\right)^{*}\mathbf{g}_{k_{j}}^{n}.$
      Adaptively select the step size $\alpha^{n}$.
$\alpha^{n}=\sigma^{2}\frac{\|\tilde{g}^{n}\|^{2}}{\sum_{j=1}^{J}\|\mathbf{H}_{k_{j}}\Pi_{M,K}^{k_{j}}\tilde{g}^{n}\|^{2}}.$
      Apply proximal update in Eqn. (18)
$f^{n+1}\leftarrow\operatorname{Prox}^{\mathcal{M}_{\Phi_{\boldsymbol{\xi}}}}_{\alpha^{n}\mathcal{R}}(f^{n}-\alpha^{n}\tilde{g}^{n})$ (23)
@@
Algorithm 1 ProxNF: Spatiotemporal image reconstruction with neural fields via proximal splitting
III-B Convergence of ProxNF
This section provides an insight on the convergence properties of Algorithm 1 in the general framework of online PnP algorithms[47] by use of well-established mathematical tools from traditional stochastic optimization and monotone operator theory.
Consider the minimization problem
$f^{*}=\operatornamewithlimits{argmin}_{f\in\mathcal{M}}{\sum_{k=1}^{K}\mathcal{L}_{k}(f)+\mathcal{R}(f)},$ (24)
where
  • A.1
    The data fidelity terms $\mathcal{L}_{k}(\cdot)$ are convex, differentiable functions with Lipschitz continuous gradients and Lipschitz constant $C>0$;
  • A.2
    The regularization term $\mathcal{R}$ is a closed proper function;
  • A.3
    The set $\mathcal{M}\subset\mathbb{U}$ is convex.
Define the operator
$P(f)\coloneqq{\rm prox}^{\mathcal{M}}_{\alpha\mathcal{R}}\left(f-\alpha\frac{K}{J}\sum_{j=1}^{J}\nabla_{f}\mathcal{L}_{k_{j}}(f)\right),$ (25)
where the operator ${\rm prox}_{\alpha\mathcal{R}}^{\mathcal{M}}$ is defined analogously to Eq.(18) and $\{k_{1},\ldots,k_{J}\}$ (with $J\ll K$) is a collection of randomly selected frames.
Under the assumptions above, Proposition 1 establishes the convergence of Algorithm 1 to a fixed point of the operator $P$, when the non-convex manifold $\mathcal{M}_{\Phi_{\boldsymbol{\xi}}}$ is replaced by the convex set $\mathcal{M}$.
Proposition 1.
Under Assumptions A.1-A.3 and with a proper choice of the step size $\alpha\in(0,1/C]$, the iteration $f^{n+1}=P(f^{n})$ converges in expectation to a fixed point of $P$, and thus to a minimizer of Eq. (24), at worse sublinearly.
Proof.
Under assumptions [A.2] and [A.3], the operator ${\rm prox}_{\alpha\mathcal{R}}^{\mathcal{M}}$ is firmly non-expansive [45], and thus $\theta$-averaged with $\theta=1/2$. Furthermore, the gradient estimate $\tilde{g}=\frac{K}{J}\sum_{j=1}^{J}\nabla_{f}\mathcal{L}_{k_{j}}(f)$ is unbiased and has bounded variance. Therefore, all the conditions of Assumption 2 in [40] hold. The convergence of the fixed-point iteration $f^{n+1}=P(f^{n})$ then follows from Proposition 5 and Corollary 1 in [40]. The equivalence between the set of the fixed points of $P$ and the minimizers of Eq. (24) follows from A.1–A.3 and the definition of the proximal operator [40]. ∎
Remark 2.
Proposition 1 provides theoretical convergence guarantees of Algorithm 1 when the NF manifold $\mathcal{M}_{\Phi_{\boldsymbol{\xi}}}$ is a convex set. While not true for a general NF architecture, this holds for the POUnet architecture in Eq. (15) with frozen partitioning parameters $\boldsymbol{\eta}$. Therefore, convergence to a fixed point can be theoretically guaranteed by modifying the proximal operator in Eq. (23), to only update the coefficients $\boldsymbol{C}$ keeping $\boldsymbol{\eta}$ fixed. Empirical evidence of the convergence in the general case is provided in the numerical results, where it was observed that also the coefficients $\boldsymbol{\eta}$ quickly stabilize after a few ProxNF iterations.
Remark 3.
Nesterov acceleration is not used Algorithm 1. This is because proximal splitting methods combining randomized evaluation of the data-fidelity gradient and/or ordered subsets with Nesterov acceleration may be susceptible to instabilities due to accumulation of errors in the momentum term [48, 49, 50, 51].
III-C Adapted partition of unity neural fields
This work also introduces a modified partition of unity network of the form
$\Phi_{(\boldsymbol{\eta},\boldsymbol{C})}(\boldsymbol{r},t)=\boldsymbol{\Psi}^{T}_{\boldsymbol{\eta}}(t)^{T}\boldsymbol{C}^{T}\boldsymbol{B}(\boldsymbol{r}),$ (26)
where the partition of unity network $\boldsymbol{\Psi}_{\boldsymbol{\eta}}$ only depends on the time coordinate $t$ and the basis functions in $\boldsymbol{B}$ only depends on the spatial coordinate $\boldsymbol{r}$. In particular, $\boldsymbol{B}(\boldsymbol{r})=\begin{pmatrix}\beta_{1}(\boldsymbol{r})&\ldots&\beta_{M}(\boldsymbol{r})\end{pmatrix}$ shares the same basis functions $\{\beta_{m}\}_{m=1}^{M}$ in Eqn. (6). This representation allows for the high spatial resolution of conventional discretization methods and smooth temporal transitions while exploiting spatiotemporal redundancies. Moreover, this architecture greatly simplifies discrete projection operations. That is for a selected frame $k$,